A 500-N tightrope walker stands at the center of the rope. If the rope can withstand a tension of 1 800 N without breaking, what is the minimum angle the rope can make with the horizontal?
1
Expert's answer
2019-10-09T10:38:11-0400

Let "T" be the tension in the rope.

"\theta" be the angle the rope make wrt. horizontal.

Vertical component of Tension will balance the weight of tight rope Walker,

So,


"2T\sin\theta=weight=500"

Now,

Maximum tension allowed = 1800 N

Therefore,


"2×1800×\sin\theta=500\\or\\\sin\theta=\frac{5}{36}\\or,\\\theta=7.983\degree"

So minimum angle that the rope can make is "7.983\degree"